# Improving Zero-shot LLM Re-Ranker with Risk Minimization

### v1

@@ Abstract
Xiaowei Yuan1,2,3, Zhao Yang1,2, Yequan Wang3, Jun Zhao1,2, Kang Liu1,2,4 1The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Beijing Academy of Artificial Intelligence 4Shanghai Artificial Intelligence Laboratory yuanxiaowei2022@ia.ac.cn, {zhao.yang, jzhao, kliu}@nlpr.ia.ac.cn tshwangyequan@gmail.com
@@ 1 Introduction
Most previous approaches trained the re-ranker on manual supervision signals with complex architectures Karpukhin et al. (2020); Nogueira et al. (2020); Formal et al. (2021), which require significant human efforts and demonstrate weak generalizability Nguyen et al. (2016); Izacard et al. (2021). As the size of models scales up (e.g., exceeding 10 billion parameters), it becomes increasingly difficult to fine-tune the dedicated re-ranking models. To address this challenge, recent efforts have attempted to leverage the zero-shot language understanding and generation capabilities of LLMs to directly enhance document re-ranking in an unsupervised way.
@@ 4.2.1 Overall Performance
Why does RankGPT performance poorly?
@@ C.1 Probability of Query Generation
According to Lafferty and Zhai (2001), when the $\hat{\theta}_{\mathbf{q}}$ is considered as the empirical distribution of the query $\mathbf{q}=q_{1}q_{2}\ldots q_{m}$; that is,
@@ Appendix E Performance on BEIR Benchmark
As illustrated in Table 5 and 6, the results demonstrate that $\mathrm{UR^{3}}$ enhances the overall rankings of the Top-100 documents both on the Mistral and GPT-Neo models. The improvements of our method are observed across all nDCG@K metrics, indicating that $\mathrm{UR^{3}}$ prioritizes relevant documents more effectively compared to the UPR method.
@@ Appendix H Discussion on Hyperparameters
Appendix I More Experiments on QA Task

### v2

@@ Abstract
Xiaowei Yuan1,2,3, Zhao Yang1,2, Yequan Wang3,∗, Jun Zhao1,2, Kang Liu1,2,∗ 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Beijing Academy of Artificial Intelligence, Beijing, China yuanxiaowei2022@ia.ac.cn, {zhao.yang, jzhao, kliu}@nlpr.ia.ac.cn tshwangyequan@gmail.com
11footnotetext: Corresponding authors.
@@ 1 Introduction
Most previous approaches trained the re-ranker on manual supervision signals Karpukhin et al. (2020); Nogueira et al. (2020); Formal et al. (2021), which require significant human efforts and demonstrate weak generalizability Izacard et al. (2021); Mokrii et al. (2021). As the size of models scales up (e.g., exceeding 10 billion parameters), it becomes increasingly difficult to fine-tune the dedicated re-ranking models. To address this challenge, recent efforts have attempted to leverage the zero-shot language understanding and generation capabilities of LLMs to directly enhance document re-ranking in an unsupervised way.
@@ 4.2.1 Overall Performance
Why does RankGPT perform poorly?
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Science and Technology Major Project (No. 2022ZD0116300) and the National Science Foundation of China (No. 62106249).
@@ C.1 Probability of Query Generation
Following the work of Lafferty and Zhai (2001), when the $\hat{\theta}_{\mathbf{q}}$ is considered as the empirical distribution of the query $\mathbf{q}=q_{1}q_{2}\ldots q_{m}$; that is,
@@
$\propto-(\sum_{w\in\mathbf{q}\cap D}\log p(w|\theta^{\prime}_{D})$
$+\sum_{w\in\mathbf{q},w\notin D}\log p(w|\theta^{\prime}_{D}))+c_{q}$
$\propto-\sum_{w\in\mathbf{q}}\log p(w|\theta^{\prime}_{D})+c_{q}$
@@ Appendix E Performance on BEIR Benchmark
As illustrated in Table 9 and 10, the results demonstrate that $\mathrm{UR^{3}}$ enhances the overall rankings of the Top-100 documents both on the Mistral and GPT-Neo models. The improvements of our method are observed across all nDCG@K metrics, indicating that $\mathrm{UR^{3}}$ prioritizes relevant documents more effectively compared to the UPR method.
@@ Appendix H Discussion on Hyperparameters
Appendix I Performance on Paraphrased Query
To evaluate the retrieval outcomes on paraphrased queries, we utilized the GPT-3.5 Turbo model to paraphrase the queries from the WebQ dataset. The paraphrasing process was guided by the prompt, "Please paraphrase the following question: question". Below are examples of the paraphrased questions we generated:
  • Original Question: "What happened after Mr. Sugihara died?"-> Paraphrased: "What occurred following Mr. Sugihara’s passing?"
  • Original Question: "Where was George Washington Carver from?"-> Paraphrased: "What was George Washington Carver’s place of origin?"
  • Original Question: "Who was Richard Nixon married to?"-> Paraphrased: "Who was the spouse of Richard Nixon?"
We then assessed the re-ranking results across various retrievers, including Contriever, BM25, MSS, and DPR, utilizing the paraphrased queries.
Llama2-7b Contriever UPR $UR^{3}$
Metric w/o para. w/ para. w/o para. w/ para.
top1 19.98 26.62 28.44 32.53 33.71
top5 43.45 54.92 56.35 58.71 59.30
top20 65.70 72.69 71.75(↓) 73.43 73.61
ndcg@1 19.98 26.62 28.44 32.53 33.71
ndcg@5 18.64 26.78 28.27 30.82 31.51
ndcg@20 22.22 31.18 32.24 33.79 34.61
MAP@100 18.79 25.92 27.12 27.82 28.67
Table 5: Performance comparison on Contriever
Llama2-7b BM25 UPR $UR^{3}$
Metric w/o para. w/ para. w/o para. w/ para.
top1 18.90 27.56 28.25 33.91 34.10
top5 41.83 54.13 54.28 55.17 56.35
top20 62.40 68.5 69.05 69.54 69.98
ndcg@1 18.90 27.56 28.25 33.91 34.10
ndcg@5 19.36 27.39 28.26 30.72 31.34
ndcg@20 22.12 31.44 32.18 33.62 34.07
MAP@100 19.15 26.63 27.18 28.09 28.45
Table 6: Performance comparison on BM25
Llama2-7b MSS UPR $UR^{3}$
Metric w/o para. w/ para. w/o para. w/ para.
top1 11.66 26.38 26.03(↓) 29.38 29.72
top5 29.04 48.67 50.34 49.85 51.48
top20 49.21 63.19 62.80(↓) 62.40 62.80
ndcg@1 11.66 26.38 26.03 29.38 29.72
ndcg@5 11.57 26.67 27.47 28.21 29.04
ndcg@20 14.84 32.46 32.77 33.20 33.61
MAP@100 12.03 26.20 26.67 26.84 27.14
Table 7: Performance comparison on MSS
Llama2-7b DPR UPR $UR^{3}$
Metric w/o para. w/ para. w/o para. w/ para.
top1 44.83 39.32 37.89(↓) 42.18 42.86
top5 65.01 66.83 64.76(↓) 66.88 66.91
top20 74.61 76.67 76.53(↓) 76.96 76.93
ndcg@1 44.83 39.32 37.89(↓) 42.18 42.86
ndcg@5 39.76 38.66 37.95(↓) 40.34 40.97
ndcg@20 38.95 41.81 41.93 42.65 43.10
MAP@100 33.32 36.46 36.63 36.82 37.44
Table 8: Performance comparison on DPR
We observed that the queries paraphrased using ChatGPT generally resulted in marginal improvements in the reranking of retrieval results both on UPR and $\mathrm{UR^{3}}$ for Contriever, BM25, and MSS. Our analysis suggests that this outcome may stem from the lower accuracy of these retrieval results to begin with. Furthermore, the distribution of queries generated by ChatGPT aligns more closely with the model data distribution than the empirical distribution of the original queries. This alignment potentially offers beneficial support for reranking initially poor retrieval outcomes.
However, for supervised retrieval systems like DPR, which already achieve high accuracy, the paraphrased queries deviate from the empirical data distribution, leading to greater negative impacts. Consequently, the performance of the UPR method is noticeably affected.
Nonetheless, our approach involves a bias-corrected estimation, which, by leveraging the document probability values, mitigates the performance decline observed in DPR results and even achieves slight improvements.
Appendix J More Experiments on QA Task